A parabola has the following characteristics:
zeros = -4, 1
axis = -1.5
V (-1.5,-2)
What is the equation for the factored form of the quadratic relation?
Please clarify.

Answer by stanbon(75887)

A parabola has the following characteristics:
zeros = -4, 1
axis = -1.5
V (-1.5,-2)
What is the equation for the factored form of the quadratic relation?
Form:
Using zeroes: y = a(x+4)(x-1)
Using V: -2 = a(-1.5+4)(-1.5-1)
-2= a(2.5)(-2.5)
a = 0.32
----
Equation:
y = 0.32(x+4)(x-1)
==========================
Cheers,
Stan H.